# Energy Landscapes for the Quantum Approximate Optimisation Algorithm

### v1

@@ II Methodology
The mixer layer performs a time-evolution of the mixer Hamiltonian $\hat{H}_{M}=-\sum_{i=1}^{N}{X_{i}}$, which anticommutes with $\hat{H}_{C}$ and is an eigenfunction of |$\ket{\psi_{0}}$. The mixer layer can be realised as a parallelisation of single-qubit parameterised $R_{x}$ quantum gates:
@@
Figure 2: (a) (left) Contour plot of $\langle\hat{H}_{C}\rangle$ for graph $G_{1}$ and its QAOA ansatz with $L=1$ against the parameters $\gamma$ and $\delta$. The solid black line depicts the minimum of each vertical slice of the contour plot, with the global minimum (GM, in green), local minimum (LM, in red) and the transition state connecting them (TS, in black) situated on the pathway that is also plotted. (b) (right) Corresponding disconnectivity graph of the contour plot in Fig. 2a, with both the GM and LM coloured based on their respective probabilities of obtaining the correct Max-Cut solution of |$\ket{\alpha\beta\alpha\beta}$.
@@ III.1 Complete unweighted graphs
Figure 5: Scatter plots of probabilities of the correct Max-Cut solutions against $\langle\hat{H}_{C}\rangle$ for individual minima of graphs $K_{6}$, $K_{7}$ and $K_{8}$ for $L=1$ (in black), $L=2$ (in red) and $L=3$ (in green).
@@
Figure 7: Scatter plots of probabilities of the correct Max-Cut solutions against $\langle\hat{H}_{C}\rangle$ for connected minima of graphs 6a8e for $L=1$ (in black), $L=2$ (in red), $L=3$ (in green) and $L=4$ (in blue).
@@ III.3 Competing QAOA Max-Cut solutions
Figure 9: Scatter plots of probabilities of competing Max-Cut solutions |$\ket{\alpha\beta\alpha\beta}$ (in red) and |$\ket{\alpha\alpha\beta\beta}$ (in blue) against $\langle\hat{H}_{C}\rangle$ for connected minima of graphs $G_{3}$, $G_{4}$ and $G_{5}$ of varying $L$.
@@ IV Conclusion
Figure 10: Disconnectivity graphs for 6a with varying circuit depth $L$.
@@
Figure 12: Disconnectivity graphs for 8a with varying circuit depth $L$.
@@
Figure 14: Disconnectivity graphs for 8c with varying circuit depth $L$.

### v2

@@ II Methodology
The mixer layer performs a time-evolution of the mixer Hamiltonian $\hat{H}_{M}=-\sum_{i=1}^{N}{X_{i}}$, which anticommutes with $\hat{H}_{C}$ and has |$\ket{\psi_{0}}$ as an eigenvector. The mixer layer can be realised as a parallelisation of single-qubit parameterised $R_{x}$ quantum gates:
@@
Figure 2: (a) (left) Contour plot of $\langle\hat{H}_{C}\rangle$ for graph $G_{1}$ and its QAOA ansatz with $L=1$ against the parameters $\gamma$ and $\delta$. The solid line depicts the minimum of each vertical slice of the contour plot, with the global minimum (GM, green circle), local minimum (LM, red triangle) and the transition state connecting them (TS, black cross) situated on the pathway that is also plotted. (b) (right) Corresponding disconnectivity graph of the contour plot in Fig. 2a, with both the GM and LM coloured (grayscaled) based on their respective probabilities of obtaining the correct Max-Cut solution of |$\ket{\alpha\beta\alpha\beta}$.
@@ III.1 Complete unweighted graphs
Figure 5: Scatter plots of probabilities of the correct Max-Cut solutions against $\langle\hat{H}_{C}\rangle$ for individual minima of graphs $K_{6}$, $K_{7}$ and $K_{8}$ for $L=1$ (black circle), $L=2$ (red square) and $L=3$ (green diamond).
@@
Figure 7: Scatter plots of probabilities of the correct Max-Cut solutions against $\langle\hat{H}_{C}\rangle$ for connected minima of graphs 6a8e for $L=1$ (black circle), $L=2$ (red square), $L=3$ (green diamond) and $L=4$ (blue triangle).
@@ III.3 Competing QAOA Max-Cut solutions
Figure 9: Scatter plots of probabilities of competing Max-Cut solutions |$\ket{\alpha\beta\alpha\beta}$ (red square) and |$\ket{\alpha\alpha\beta\beta}$ (blue circle) against $\langle\hat{H}_{C}\rangle$ for connected minima of graphs $G_{3}$, $G_{4}$ and $G_{5}$ of varying $L$.